# Massive $\nu$s through the CNN lens: interpreting the field-level neutrino mass information in weak lensing

### v2

@@ Acknowledgments
We thank Vanessa Böhm, Fraçoise Lanusse, Jia Liu, Uroš Seljak, Francisco Villaescusa-Navarro, and Benjamin Wandelt for insightful discussion. MG also thanks Vanessa Böhm for her initial mentorship. This research was conducted as part of the N3AS Undergraduate Research Program at UC Berkeley, supported by the NSF Award Number 2020275 and Heising Simons Grant 2017-228. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award ASCR-ERCAP0029232. The computations reported in this paper were (in part) performed using resources made available by the Flatiron Institute. The analysis of the simulations has made use of the Pylians library, publicly available at https://github.com/franciscovillaescusa/Pylians3.

### v3

@@ Acknowledgments
We thank Vanessa Böhm, Françoise Lanusse, Jia Liu, Uroš Seljak, Francisco Villaescusa-Navarro, and Benjamin Wandelt for insightful discussion. MG also thanks Vanessa Böhm for her initial mentorship. This research was conducted as part of the N3AS Undergraduate Research Program at UC Berkeley, supported by the NSF Award Number 2020275 and Heising Simons Grant 2017-228. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award ASCR-ERCAP0029232. The computations reported in this paper were (in part) performed using resources made available by the Flatiron Institute. The analysis of the simulations has made use of the Pylians library, publicly available at https://github.com/franciscovillaescusa/Pylians3.